All systems are go

Don't agree with this? Please let us know

  • I-DASH WEBSITE
  • CUSTOMER SUPPORT PORTAL
  • REMOTE ACCESS PORTAL
  • HOSTED SERVICES
    • HOSTED SERVERS
    • CLOUDMANAGED NETWORKING
    • WEB HOSTING
  • HORIZON CLOUD PBX
  • NETWORK INFRASTRUCTURE
    • I-DASH DC01 (MILTON KEYNES)
    • I-DASH DC02 (ST NEOTS)
Previous Incidents

[Resolved] Horizon Telephone Systems - Gamma Network

This incident lasted 8 hours and 31 minutes.
Wed, 14 Nov 2018
10:40:11 GMT

Our support team is currently aware of a nationwide issue affecting the Gamma Network, this issue is also affecting some Horizon Telephone systems.
Whilst this is out of our control, we are working to support the provider in order to rectify the situation as soon as possible.
our i-dash support team will update you further as and when we have useful information to share.

11:15:42 GMT

An emergency patch was applied to the impacted application cluster and we have monitored a positive recovery, it is now replicating this work across remaining application clusters, to ensure full stability

11:16:20 GMT

We've fixed the core issue, and are waiting for things to recover.

11:32:13 GMT

Our engineering teams are continuing to work with our vendor whilst they replicate the earlier work across the remaining clusters to ensure full stability.
A further update will be issued within the next 30 minutes. We apologise for the ongoing impact to service, this incident is being managed as our highest priority.

12:24:51 GMT

We are continuing to see some intermittent issues across our Horizon base and the Supplier are continuing to work on full restore as a priority.

13:07:17 GMT

We are seeing early signs of recovery following a patch of the root cause and further stabilisation measures which have been implemented.
We do not have a current indication of when all services will be fully restored, but we are taking further steps to ensure this happens as quickly as possible. We continue to monitor the situation and will advise on the recovery progress within the next 30 minutes.

14:13:40 GMT

We are still monitoring a slow but positive recovery of Horizon services following the earlier work.

We are now taking some additional measures with our technology partner to expedite the recovery.

Whilst the recovery is underway, customers may still experience partial or no service on the Horizon platform currently.

14:24:43 GMT

We have now taken further action with our technology partner to further stabilise the Horizon service and we expect to start to see an improvement on the ability for customers to make and receive calls consistently by 15:00. We have also increased performance on the Horizon portal, which should increase availability of the portal for customers during the incident.

We apologise for the ongoing disruption to service.

15:31:51 GMT

After performing some changes on core equipment we have seen a significant uplift in the number of handsets successfully registering from approximately 14:50. The issue is ongoing and a number of customers remain affected, we are continuing to work on the incident as a priority. There are additional works planned over the next 30 minutes that should help to further improve the registration. On newly registered handsets we have had a small number or reports of one way audio, or no audio, which we are also investigating as a priority. We will provide a further update in 30 minutes.

15:46:14 GMT

The Supplier have now taken further action with our technology partner to further stabilise the Horizon service and we expect to start to see an improvement on the ability for customers to make and receive calls consistently. They have also increased performance on the Horizon portal, which should increase availability of the portal for customers during the incident.

We apologise for the ongoing disruption to service.

16:02:40 GMT

There are a number of customers that remain affected by the registration issues, although we have seen the issue resolved for some, this is not the case for all customers. Our engineering teams continue to work closely with our technology partners attempting a number of changes to work towards a full restoration. We apologise for the ongoing issues

16:33:21 GMT

The Supplier is reporting that they continue to prioritise this incident and work toward full restoration for those customers that remain affected. We apologise for the ongoing disruption

16:44:39 GMT

The issues with registration continues to improve and is a good indicator that the Horizon platform is stabilising, although not completely resolved for all customers.
Some registered handsets are not fully operational and although now registered could be unresponsive, or have audio issues on calls. Our teams continue to investigate all options to restore full service to all customers.

18:33:18 GMT

We have completed remaining changes to the Horizon network and we have seen normal service and traffic levels resume as of 17:48. For any customers with ongoing issues we kindly request they reboot their Horizon handsets, for any ongoing issues please contact the relevant support team for further diagnostics. There are reports of Horizon Portal pages slow to load or access, and this is due to an influx of requests to remove diverts placed earlier and we expect portal performance to improve again shortly. We will be planning some further works tonight and will provide an update as to the planned timings within the next hour. We sincerely apologies for the impact to all affected customers. The next update will be at 7PM.

19:11:41 GMT

As per our last update, we are planning some additional work on our Horizon platform this evening. The work is to bring the platform in to normal operating parameters following some earlier tactical configuration changes made. We do not anticipate that this work will further disrupt Horizon services, but they will be at risk until the changes are completed. We will continue to monitor the network throughout the evening but remain confident that we have fully resolved the earlier issues and will have a normal working service for customers into tomorrow morning.

[Resolved] Service Outage Detected

This incident lasted 3 hours and 51 minutes.
  • CUSTOMER SUPPORT PORTAL
Thu, 18 Oct 2018
05:18:17 BST

Our monitoring has detected that we may be experiencing issues with this service, Our engineers have been made aware and are looking into the issue.

09:09:51 BST

Services have been restored and are confirmed as online via our monitoring.

[Complete] Scheduled Maintenance - Gw6 MKDC-0

This was a planned incident, it took 29 minutes.
  • HOSTED SERVERS
  • CLOUDMANAGED NETWORKING
  • I-DASH DC01 (MILTON KEYNES)
Wed, 10 Oct 2018
11:08:52 BST
Overview

Network engineers will be performing maintenance on gw6.rtr.mkdc0.as45014.net at MKDC-0 These works are to perform an EOS upgrade:-

• Upgrade to EOS 4.21

- Reason for works: Upgrade EOS image to latest 4.21 train to add additional MPLS features

  • Scheduled for Tue, 23 Oct 2018 8:00PM BST
  • Expected to take 30 minutes
Tue, 23 Oct 2018
20:00:03 BST

The scheduled maintenance is now underway. We'll keep you updated on our progress.

20:30:01 BST

The maintenance is now complete. Thanks for your patience.

[Resolved] Connectivity Issue affecting Milton Keynes Network

This incident lasted 1 day and 7 minutes.
  • REMOTE ACCESS PORTAL
  • HOSTED SERVICES
  • HOSTED SERVERS
  • CLOUDMANAGED NETWORKING
  • WEB HOSTING
  • I-DASH DC01 (MILTON KEYNES)
Tue, 25 Sep 2018
16:39:47 BST

Engineers are aware of an issue affecting customers on our Milton Keynes network. This is currently being investigated and further updates will follow.

16:56:54 BST

Engineers have identified a possible fault with Core02 at MKDC.
An emergency reload of Core02 will take place.

19:06:41 BST

Service is currently stable and engineers are continuing to monitor Core02 and the overall network before marking this as resolved in case any further issues develop.

23:42:40 BST

Engineers believe the fault in Core02 to be a hardware component issue and are organising for this component to be replaced. In the meantime, all traffic continues to use Core01 and network status will continue to be monitored.

Wed, 26 Sep 2018
16:47:10 BST

Engineers will continue to monitor closely however are now satisfied that this has now been resolved and Core02 has been stable for 9 hours.

[Complete] VMware Host Patching

This was a planned incident, it took 2 hours.
  • REMOTE ACCESS PORTAL
  • HOSTED SERVERS
  • CLOUDMANAGED NETWORKING
Fri, 24 Aug 2018
13:26:03 BST
Overview

Patches and rolling reboots will be applied to hosts in the cluster.
Performance may be effected - Services will remain online.

  • Scheduled for Fri, 24 Aug 2018 1:30PM BST
  • Expected to take 2 hours
13:30:03 BST

The scheduled maintenance is now underway. We'll keep you updated on our progress.

15:30:02 BST

The maintenance is now complete. Thanks for your patience.

[Complete] Web Hosting Farm Patching

This was a planned incident, it took 1 hour and 3 seconds.
  • CUSTOMER SUPPORT PORTAL
  • WEB HOSTING
Thu, 23 Aug 2018
19:49:53 BST
Overview

Websites hosted on the web hosting farm will be intermittently unavailable while patches are applied and rolling reboots issued.

  • Scheduled for Thu, 23 Aug 2018 9:00PM BST
  • Expected to take 1 hour
21:00:01 BST

The scheduled maintenance is now underway. We'll keep you updated on our progress.

22:00:03 BST

The maintenance is now complete. Thanks for your patience.

No further notices from the past 90 days.